# Can We Catch the Elephant? The Evolvement of Hallucination Evaluation on Natural Language Generation: A Survey

### v1

@@ 5 Conclusion and Future Direction
Evaluating hallucination in NLG is crucial as it determines the direction and trend for developing a better model. Historically, most metrics are task-based, as defining the criteria for task performance is relatively straightforward. As NLP research gradually shifts to LLMs, researchers have proposed new demands and challenges. We summarize several aspects of hallucination evaluation that urgently need to be addressed in the future.
Comprehensive evaluation Traditional evaluation methods based on a specific task are no longer sufficient to LLMs. Researchers have already proposed many comprehensive evaluation dimensions for the general capabilities, such as common sense and reasoning. A comprehensive and unified benchmark specialized for hallucination in the future will greatly promote model iteration to address hallucination issues.
@@
Differentiation between hallucination and error Hallucinations and text errors can be confusing sometimes. According to the traditional definition of hallucination (smooth but incorrect), any response from a large model that differs from the ground truth can be considered as hallucination, which is obviously unreasonable and can mislead researchers. Some works in NMT have already made some progress (Dale et al., 2023), and it’s better for future evaluation methods to accurately identify real hallucinations.
@@
Interpretabity Identifying fact granularity and analyzing hallucination reasons can provide significant assistance in solving hallucination problems. Some reasoning methods have the potential to analyze the underlying causes of hallucinations and have better evaluation. Other observing aspects lie in the internal state of model generation (Su et al., 2024), model structures, and even data distributions, which provide more analytical perspectives.
@@
Long context/generation Current exploration of LLMs in long-context scenarios mainly focuses on long inputs, such as using a complete contract or multi-turn dialogue history as input. It is crucial to address hallucinations caused by the model’s difficulty in understanding long inputs. However, on the other hand, hallucinations of the type caused by contradictions between the beginning and end of long outputs are also worth exploring (Wei et al., 2024), such as detecting whether a model-generated novel contains inconsistencies in character behavior.
@@
More application views LLMs have more challenging hallucinations in certain scenarios. In the multi-lingual scenario, low-resource languages other than English may be more prone to hallucinations, and cross-lingual conversation can face misunderstanding hallucinations due to different cultural backgrounds. Moreover, the latest research focuses on expanding LLMs to areas such as autonomous agents, multimodality, and real-world applications, which bring about new types of hallucinations, such as code hallucination (Qian et al., 2023), tool hallucination, visual hallucination, and so on. Detecting such hallucinations is very interesting and worthwhile direction to explore.
@@ Appendix C Supplement Table
Era Name New Dataset Data Source Fact Definition Task Based-model Method Metric SF WF Before LLM with Ref $Fact_{acc}$ WikiFact Wikipedia, Wikidata KB Ent-rel -ent Summ Transformer NER, Binary Relation Classifier P, R, F1 FactCC FactCC CNN/DM, XSumFaith Sent Summ BERT NLI (2-class) Likelihood DAE DAE PARANMT50M Dependency Summ ELECTRA NLI (2-class) Likelihood Maskeval / CNN/DM, WikiLarge, ASSET Word Summ, Simp T5 Word Weighting Weighted Match Score  Guerreiro et al. (2023) Haystack WMT2018 DE-EN Text Span NMT Transformer Uncertainty Measure Avg. Similarity  Dale et al. (2023) / Haystack Text Span NMT Transformer Source Contribution Percentage Before LLM Ref Free FEQA FEQA CNN/DM, XSum Sent Span Summ BART (QG), BERT (QA) QG-QA Avg. F1 QAGS QAGS CNN/DM, XSum Ent, Noun Phrase Summ BART (QG), BERT (QA) QG-QA Avg. Similarity QuestEval / CNN/DM, Xsum Ent, Noun Summ T5 (QG, QA) QG-QA P, R, F1 QAFactEval / SummaC NP Chunk Summ BART (QG), ELECTRA (QA) QG-QA, NLI LERC MQAG / QAGS, XSumFaith, Podcast Assessment, SummEval Sent Span Summ T5 (QG), Longformer (QA) Multi-Choice QA Choice Statistical Distance CoCo / QAGS, SummEval Token, Span, Sent, Doc Summarization BART Counterfactual Estimation Avg. Likelihood Diff ExtEval ExtEval CNN/DM Discourse, Coreference, Sentiment Summ SpanBERT, RoBERTa Direct Prediction, Statistic Summation of Sub-scores $Q^{2}$ $Q^{2}$ WOW Sent Span Diag T5 (QG); Albert- Xlarge (QA); RoBERTa (NLI) QG-QA, NLI Likelihood AlignScore / 22 datasets from 7 tasks Sent NLI, QA, paraphrasing, Fact Verification, IR, Semantic Similarity, Summ RoberTa 3-way classification Likelihood WeCheck / TRUE Response Summ, Diag, Para, Fact Check DeBERTaV3 Weakly Supervised NLI Likelihood After LLM tool SCALE ScreenEval LLM, Human Sentence Long Diag Flan-T5 NLI Likelihood  Chen et al. (2023b) / SummEval, XSumFaith, Goyal21, CLIFF Response Summarization Flan-T5, code-davinci-002, text-davinci-003, ChatGPT, GPT-4 Vanilla/COT/ Sent-by-Sent Prompt Balanced Acc GPTScore / 37 datasets from 4 tasks Various Summ, Diag, NMT, D2T GPT-2, OPT, FLAN, GPT-3 Direct Assessment Direct Score G-Eval / SummEval, Topical-Chat, QAGS Response Summ, Diag GPT-4 COT, Form-filling Weighted Scores Wang et al. (2023b) / 5 datasets from 3 tasks Response Summ, D2T, Story Gen ChatGPT Direct Assessment, Rating Direct score ChainPoll RealHall-closed, RealHall-open COVID-QA, DROP, Open Ass prompts, TriviaQA Response Hallu Detect gpt-3.5-turbo Direct Assessment (2-class) Acc After LLM object FacTool / RoSE, FactPrompts, HumanEval, GSM-Hard, Self-instruct Claim, Response Knowledge-based QA, Code Generation, Math Reasoning, Sci-literature Review ChatGPT Claim Extraction, Query Generation, Tool Querying, Evidence Collection, Agreement Verification P, R, F1 UFO / NQ, HotpotQA, TruthfulQA, CNN/DM, Multi-News, MS MARCO Ent Open-domain/ Web Retrieval-based/ Expert-validated/ Retrieval-Augmented QA, News Fact Generation ChatGPT (gpt-3.5-turbo-1106) Fact Unit Extraction, Fact Source Verification, Fact Consistency Discrimination Avg. Sub-scores CONNER / NQ, WoW Sentence Open-domain QA, Knowledge-grounded Dialogue NLI-RoBERTa -large, ColBERTv2 3-way NLI Acc SelfCheckGPT SelfCheckGPT WikiBio Response Hallu Detect GPT-3 Direct Assessment, QA, BERTScore, NLI, Ngram AUC-PR $SAC^{3}$ / HotpotQA, NQ-open Response QA Generation gpt-3.5-turbo, Falcon-7b-instruct, Guanaco-33b Cross-checking, QA Pair Consistency AUROC EigenScore / CoQA, SQuAD, TriviaQA Natural Questions Inner State Open-book QA Closed-book QA LLaMA, OPT Semantic Consistency/ Diversity in Dense Embedding Space AUROC, PCC TruthfulQA TruthfulQA LLM, Human Response Multi-Choice QA, Generation GPT-3-175B Answer Match Score HaluEval Task-specific, General Alpaca, Task datasets ChatGPT Response QA, Summ, Knowledge- grounded Diag, Generation ChatGPT Direct Assessment Acc FACTOR Wiki-/News-/ Expert- FACTOR Wikipedia, Refin- edWeb, ExpertQA Sent Span Generation / FRANK Error Classification likelihood FELM FELM TruthfulQA, Quora, MMLU, GSM8K, ChatGPT, Human Text Span, Claim World Knowledge, Sci and Tech, Math, Writing and Recom- mendation, Reasoning Vicuna, ChatGPT, GPT4 Direct Assessment F1, Balanced Acc FreshQA Never/Slow Fast-changing, false-premise Human Response Generation / Answer Match Acc KoLA KoLA Wikipedia, Updated News and Novels Response Knowledge Memorization /Understanding/Applying /Creating / Self-contrast Answer Match Similarity RealTimeQA RealTimeQA CNN, THE WEEK, USA Today Response Multi-Choice QA, Generation GPT-3, T5 Answer Match Acc, EM, F1 ERBench ERBench Database 5 datasets from Kaggle Ent-Rel Binary/ Multiple -choice QA / Direct Assessment, String Matching Ans/Rat/ Ans-Rat Acc, Hallu Rate FactScore / Biographies in Wikipedia Atomic Fact Generation InstructGPT, ChatGPT, PerplexityAI Binary Classification P BAMBOO SenHallu, AbsHallu 10 datasets from 5 tasks Response Multi-choice tasks, Select tasks ChatGPT Answer Match P, R, F1 RV PHD Human Annotator Ent Generation ChatGPT Construct Query, Access Databases, Entity-Answer Match P, R, F1 SummEdits SummEdits 9 datasets from Summ task Span Summ, Reasoning gpt-3.5-turbo Seed summary verify, Summary edits, Annotation Balanced Acc MedHalt MedHalt MedMCQA, Medqa USMILE, Medqa (Taiwan), Headqa, PubMed Response Reasoning Hallu Test, Memory Hallu Test ChatGPT Answer Match Pointwise Score, Acc ChineseFactEval ChineseFactEval / Response Generation / FacTool, Human annotator Direct Score UHGEval UHGEval Chinese News Websites Keywords Generative/ Discriminative/ Selective Evaluator GPT-4 Answer Match, Similarity Acc, Similarity Score HalluQA HalluQA Human Response Generation GLM-130B, ChatGPT, GPT-4 Direct Assessment Non-hallu Rate

### v2

@@ 5 Conclusion and Future Direction
Evaluating hallucination in NLG is crucial as it determines the direction and trend for developing better models. Historically, most metrics have been task-based, as defining the criteria for task performance is relatively straightforward. However, as NLP research gradually shifts to LLMs, researchers have proposed new demands and challenges. We summarize several aspects of hallucination evaluation that urgently need to be addressed in the future.
@@
Hallucination and text error It can be challenging to distinguish between hallucinations and text errors. According to the traditional definition of hallucination (smooth but incorrect), any response from a large model that differs from the ground truth can be considered as hallucination, which is obviously unreasonable and can mislead researchers. Some works in NMT have already made progress in this area (Dale et al., 2023), and future evaluation methods should aim to accurately identify real hallucinations.
Hallucination and imagination Is hallucination always bad? Not necessarily. In certain contexts, such as chatting about a sci-fi novel, the content is rich with imagination, and the dialogue should be creative as well. In this circumstance, the distinction between hallucination and imagination is quite subtle. Distinguishing between these two phenomena can help models evaluate more diverse texts.
Supporting theories So far, the evaluation and detection of hallucination have been limited to observing the final output of the model – detecting hallucination manifested in the text. If there is theoretical support for researchers to detect hallucination from internal hidden states, model structures, and even data distributions, it can greatly reduce the cost of model iteration.
@@
Interpretability Identifying fact granularity and analyzing the reasons behind hallucination can provide significant assistance in solving hallucination problems. Some reasoning methods have the potential to analyze the underlying causes of hallucinations and offer better evaluation. Other observing aspects lie in the internal state of model generation (Su et al., 2024), which provide more analytical perspectives.
@@
Long context/generation Current exploration of LLMs in long-context scenarios mainly focuses on long inputs, such as using a complete contract or multi-turn dialogue history as input. It is crucial to address hallucinations caused by the model’s difficulty in understanding long inputs. However, hallucinations caused by contradictions between the beginning and end of long outputs are also worth exploring (Wei et al., 2024), such as detecting whether a model-generated novel contains inconsistencies in character behavior.
@@
More application views LLMs have more challenging hallucinations in certain scenarios. In the multilingual scenario, low-resource languages other than English may be more prone to hallucinations, and cross-lingual conversation can face misunderstanding hallucinations due to different cultural backgrounds. Moreover, the latest research focuses on expanding LLMs to areas such as autonomous agents, multimodality, and real-world applications, which bring about new types of hallucinations, such as code hallucination (Qian et al., 2023), tool hallucination, visual hallucination, multimodal hallucination, and so on. Detecting such hallucinations is a very interesting and worthwhile direction to explore.
6 Limination
In this paper, we collect a broad range of related papers and reports, categorized and compared the methods and offered our insights for future directions. However, this paper does have some limitations. First of all, we did not do comprehensive experiments to revisit the above evaluators, because the evaluators usually focus on different types of hallucinations for various tasks, and it wouldn’t be fair to compare across the categories. For example, evaluators for LLMs intend to build their own datasets with human annotation, which vary in categories and schemes. Second, we put the content related to fact-checking and human evaluation in Appendix AB. We do not have enough pages to provide a thorough explanation of each of them, and detailed descriptions could also blur the focus of this paper. At last but not least, the case study we provide in Appendix D only includes a few representative cases on selective models for reference.
Acknowledgments
@@ Appendix C Supplement Table
Appendix D Case Study
Here we present some results of selected evaluators on different kinds for SF or WF errors on summarization data in Table 2. The data we used are from XEnt dataset (Cao et al., 2022) and FactCollect (Ribeiro et al., 2022). We selected evaluators that use the GPT series and those that do not, covering both models that evaluate SF and WF facets. For the models utilizing LLMs, we specifically employed GPT-3.5-turbo. Although FacTool is not directly applicable for evaluating summarization tasks, we conducted experiments under its KBQA (Knowledge-Based Question Answering) setup to see its transfer ability.
The results of different models on these cases show considerable variation. In the SF-WF case, only FacTool made an incorrect judgment, which might be attributed to its insufficient transfer ability. SelfCheckGPT uses a zero-shot approach in its prompt to assess the consistency, whereas HaluEval’s prompt provides examples for judgment. However, the SFE cases indicate that the results of these two evaluators remain unstable. For the WFE cases, FacTool provides the correct answers, and surprisingly, WeCheck also made correct judgments. Currently, to our best knowledge, no such labeled data is available for full evaluation. More accurate data is needed for further experiments to validate the preferences of different evaluators.
@@ Appendix D Case Study
Era Name New Dataset Data Source Fact Definition Task Based-model Method Metric SF WF Before LLM Ref -based $Fact_{acc}$ WikiFact Wikipedia, Wikidata KB Triplet Summ Transformer NER, Classification P, R, F1 FactCC FactCC CNN/DM, XSumFaith Sent Summ BERT NLI (2-class) Likelihood DAE DAE PARANMT50M Dependency Summ ELECTRA NLI (2-class) Likelihood Maskeval / CNN/DM, WikiLarge, ASSET Word Summ, Simp T5 Word Weighting Weighted Match Score  Guerreiro et al. (2023) Haystack WMT2018, DE-EN Text Span NMT Transformer Uncertainty Measure Avg. Similarity  Dale et al. (2023) / Haystack Text Span NMT Transformer Source Contribution Percentage Before LLM Ref -free FEQA FEQA CNN/DM, XSum Sent Span Summ BART (QG), BERT (QA) QG-QA Avg. F1 QAGS QAGS CNN/DM, XSum Ent, Noun Phrase Summ BART (QG), BERT (QA) QG-QA Avg. Similarity QuestEval / CNN/DM, Xsum Ent, Noun Summ T5 (QG, QA) QG-QA P, R, F1 QAFactEval / SummaC NP Chunk Summ BART (QG), ELECTRA (QA) QG-QA, NLI LERC MQAG / QAGS, XSumFaith, Podcast, Assessment, SummEval Sent Span Summ T5 (QG), Longformer (QA) Multi-Choice QA Choice Statistical Distance CoCo / QAGS, SummEval Token, Span, Sent, Doc Summ BART Counterfactual Estimation Avg. Likelihood Diff FactGraph FactCollect CNN/DM, XSum Dependency Summ ELECTRA Classification BACC, F1 FactKB FactKB CNN/DM, XSum Triplet Summ RoBERTa Classification BACC, F1 ExtEval ExtEval CNN/DM Discourse, Coreference, Sentiment Summ SpanBERT, RoBERTa Direct Prediction, Statistic Summation of Sub-scores $Q^{2}$ $Q^{2}$ WOW Sent Span Diag T5 (QG), Albert-Xlarge (QA), RoBERTa (NLI) QG-QA, NLI Likelihood FactPush / TRUE Span Diag, Summ, Paraphrase DeBERTa NLI AUC AlignScore / 22 datasets from 7 tasks Sent NLI, QA, Paraphrase, Fact Verification, IR, Semantic Similarity, Summ RobERTa 3-way Classification Likelihood WeCheck / TRUE Response Summ, Diag, Para, Fact Check DeBERTaV3 Weakly Supervised NLI Likelihood After LLM Direct SCALE ScreenEval LLM, Human Sentence Long Diag Flan-T5 NLI Likelihood  Chen et al. (2023b) / SummEval, XSumFaith, Goyal21, CLIFF Response Summ Flan-T5, code-davinci-002, text-davinci-003, ChatGPT, GPT-4 Vanilla/COT/ Sent-by-Sent Prompt Balanced Acc GPTScore / 37 datasets from 4 tasks Various Summ, Diag, NMT, D2T GPT-2, OPT, FLAN, GPT-3 Direct Assessment Direct Score G-Eval / SummEval, Topical-Chat, QAGS Response Summ, Diag GPT-4 COT, Form-filling Weighted Scores Wang et al. (2023b) / 5 datasets from 3 tasks Response Summ, D2T, Story Gen ChatGPT Direct Assessment, Rating Direct score ChainPoll RealHall-closed, RealHall-open COVID-QA, DROP, Open Ass prompts, TriviaQA Response Hallu Detect gpt-3.5-turbo Direct Assessment (2-class) Acc EigenScore / CoQA, SQuAD, TriviaQA Natural Questions Inner State Open-book QA Closed-book QA LLaMA, OPT Semantic Consistency/ Diversity in Dense Embedding Space AUROC, PCC TruthfulQA TruthfulQA LLM, Human Response Multi-Choice QA, Generation GPT-3-175B Answer Match Score HaluEval Task-specific, General Alpaca, Task datasets ChatGPT Response QA, Summ, Knowledge- grounded Diag, Generation ChatGPT Direct Assessment Acc FACTOR Wiki-/News-/ Expert- FACTOR Wikipedia, Refin- edWeb, ExpertQA Sent Span Generation / FRANK Error Classification likelihood FELM FELM TruthfulQA, Quora, MMLU, GSM8K, ChatGPT, Human Text Span, Claim World Knowledge, Sci and Tech, Math, Writing and Recom- mendation, Reasoning Vicuna, ChatGPT, GPT4 Direct Assessment F1, Balanced Acc FreshQA Never/Slow Fast-changing, false-premise Human Response Generation / Answer Match Acc RealTimeQA RealTimeQA CNN, THE WEEK, USA Today Response Multi-Choice QA, Generation GPT-3, T5 Answer Match Acc, EM, F1 ERBench ERBench Database 5 datasets from Kaggle Ent-Rel Binary/ Multiple -choice QA / Direct Assessment, String Matching Ans/Rat/ Ans-Rat Acc, Hallu Rate FactScore / Biographies in Wikipedia Atomic Fact Generation InstructGPT, ChatGPT, PerplexityAI Binary Classification P BAMBOO SenHallu, AbsHallu 10 datasets from 5 tasks Response Multi-choice tasks, Select tasks ChatGPT Answer Match P, R, F1 MedHalt MedHalt MedMCQA, Medqa USMILE, Medqa (Taiwan), Headqa, PubMed Response Reasoning Hallu Test, Memory Hallu Test ChatGPT Answer Match Pointwise Score, Acc ChineseFactEval ChineseFactEval / Response Generation / FacTool, Human annotator Direct Score UHGEval UHGEval Chinese News Websites Keywords Generative/ Discriminative/ Selective Evaluator GPT-4 Answer Match, Similarity Acc, Similarity Score HalluQA HalluQA Human Response Generation GLM-130B, ChatGPT, GPT-4 Direct Assessment Non-hallu Rate After LLM Pipeline FacTool / RoSE, FactPrompts, HumanEval, GSM-Hard, Self-instruct Claim, Response Knowledge-based QA, Code Generation, Math Reasoning, Sci-literature Review ChatGPT Claim Extraction, Query Generation, Tool Querying, Evidence Collection, Agreement Verification P, R, F1 UFO / NQ, HotpotQA, TruthfulQA, CNN/DM, Multi-News, MS MARCO Ent Open-domain/ Web Retrieval-based/ Expert-validated/ Retrieval-Augmented QA, News Fact Generation ChatGPT (gpt-3.5-turbo-1106) Fact Unit Extraction, Fact Source Verification, Fact Consistency Discrimination Avg. Sub-scores CONNER / NQ, WoW Sentence Open-domain QA, Knowledge-grounded Dialogue NLI-RoBERTa -large, ColBERTv2 3-way NLI Acc SelfCheckGPT SelfCheckGPT WikiBio Response Hallu Detect GPT-3 NLI, Ngram, QA, BERTScore, Prompt AUC-PR $SAC^{3}$ / HotpotQA, NQ-open Response QA Generation gpt-3.5-turbo, Falcon-7b-instruct, Guanaco-33b Cross-checking, QA Pair Consistency AUROC KoLA KoLA Wikipedia, Updated News and Novels Response Knowledge Memorization /Understanding/Applying /Creating / Self-contrast Answer Match Similarity RV PHD Human Annotator Ent Generation ChatGPT Construct Query, Access Databases, Entity-Answer Match P, R, F1 SummEdits SummEdits 9 datasets from Summ task Span Summ, Reasoning gpt-3.5-turbo Seed summary verify, Summary edits, Annotation Balanced Acc
@@
Document Summary Note WeCheck SelfCheckGPT HaluEval FacTool SF-WF … Harry Kane has been given the nod by Youssouf Mulumbu for this season’s players’ Player of the Year award The West Brom midfielder has picked Chelsea wideman Eden Hazard for the young player of the year prize Congo international Mulumbu posted his votes for this year’s PFA awards to Twitter on Wednesday Mulumbu challenges QPR defender Yun Suk-Young during West Brom’s 4-1 defeat at The Hawthorns Goalkeepe … The DR Congo international has picked Chelsea wideman Eden Hazard for the young player of the year prize . The summary is correct. TRUE TRUE TRUE FALSE SF-WFE … Since the end of March, the Vikings’ only wins have been in the Challenge Cup against lower-league sides. "We’ve got the personnel and we’ve got the people to spark us back into life," Chris Betts told BBC Radio Merseyside. "When we get rolling again I’m sure, or I’m positive, that we can really turn this year around for ourselves." … "The players are hurting and we’ve got to win," added England assistant coach Betts. … Widnes Vikings can turn their poor start to the Super League season around if they can find a winning streak, says assistant coach Chris Betts. "Chris Betts" is in the document but is incorrect essentially. FALSE TRUE TRUE FALSE SFE-WF The panther chameleon was found on Monday by a dog walker in the wooded area at Marl Park. It had to be put down after X-rays showed all of its legs were broken and it had a deformed spine. RSPCA Cymru said it was an "extremely sad example of an abandoned and neglected exotic pet". …… A chameleon has been put down by RSPCA Cymru after it was found injured and abandoned in a Cardiff park. The Marl Park is in Cardiff but not mentioned in the document. TRUE FALSE TRUE TRUE SFE-WFE A number of men, two of them believed to have been carrying guns, forced their way into the property at Oakfield Drive shortly after 20:00 GMT on Saturday. They demanded money before assaulting a man aged in his 50s. … Alliance East Antrim MLA Stewart Dickson has condemned the attack. … A man has been assaulted by a gang of armed men during a robbery at a house in Ballymena, County Antrim. "Ballymena" is neither in the document nor correct according to external knowledge. FALSE TRUE TRUE FALSE
Table 2: Examples of the results from selected evaluators on the SF errors (SFE) and WF errors (WFE).